OpenStudy (anonymous):

The greatest wind speed recorded on Earth’s surface is 231 miles per hour. What is the speed in kilometers per hour?

OpenStudy (paxpolaris):

how many miles in km?

OpenStudy (anonymous):

144 372 610 1,274 those are the answer choices

OpenStudy (paxpolaris):

\[1mile \approx1.61km\]

OpenStudy (anonymous):

so im going from large to small which is....multiply right?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

the answer is 372? :]

OpenStudy (paxpolaris):

\[\Large 231{\cancel{mile}\over hr} \times {1.61km \over \cancel{mile}}=\] yes

OpenStudy (anonymous):

ohk now how do i turn it into hours?

OpenStudy (anonymous):

or was I right the first time?

OpenStudy (paxpolaris):

=372km/hr

OpenStudy (paxpolaris):

you're right

OpenStudy (anonymous):

thank you. I have another one do you mind?

OpenStudy (paxpolaris):

no go ahead

OpenStudy (anonymous):

An event lasts 1,920,000 milliseconds. Approximately how long did the event last in hours? I guess 0.53 for that is it right?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

ohk one more thing....do you have a picture of measurements?

OpenStudy (paxpolaris):

what do you mean?

OpenStudy (anonymous):

i need measurements for kilo,deci,centi's all those do you have any i could look at?

OpenStudy (paxpolaris):

kilo = 1000 times hecta= 100* deca = 10* deci = 1/10 centi= 1/ 100 milli = 1/1000 micro = 1/1,000,000 nano = 1/1,000,000,000

OpenStudy (anonymous):

:D Thanks!

OpenStudy (anonymous):

An event lasts 1,320,000 milliseconds. Approximately how long did the event last in hours? 0.37 3.7 37 370

OpenStudy (anonymous):

please help me :)